OpenStudy (anonymous):

2x-1/3=2x-1/3 the 2x-1 on both sides are on the top of the fraction and the 3's are on the bottom

OpenStudy (anonymous):

@SolomonZelman please help

OpenStudy (anonymous):

so what's the question then if: \[\frac{ 2x -1 }{ 3 } = \frac{ 2x-1 }{ 3 }\]

OpenStudy (anonymous):

dont you have to solve it im trying to tell if its a contradiction or identity @Adjax

OpenStudy (anonymous):

2x-1 / 3 = 2x-1 / 3 Multiply both sides by 3 2x-1 = 2x-1 Add 1 to both sides 2x =2x Divide both sides by 2 x=x So, this is saying that x could be any number.

jhonyy9 (jhonyy9):

so the left side equal right side